O 0.05 QUESTION 6 You have three jars containing marbles, as follows: ar 1: 600 red and 400 white jar 2:900 blue and 100 white jar 3: 10 green and 990 white If you randomly select one marble from each jar, calculate the probability of obtaining a red, a blue, and a green. 00:0054 0:09946 0 00006 0,0396 QUESTION 7 How many genotypes are there in the 16 squares of the Punnett square below?
Answer #1

Answer:

Based on the given information:

  • Jar 1 contains: 600 red and 400 white marbles, total marbles = 600 + 400 = 1000
  • Jar 2 contains: 900 blue and 100 white marbles, total marbles = 900 + 100 = 1000
  • Jar 3 contains: 10 green and 990 white marbles, total marbles = 10 + 990 = 1000
  • Probability of obtaining a red marble from first jar P(R) = Number of red marbles / Total number of marbles = 600/1000 = 0.6
  • Probability of obtaining a blue marble from second jar P(B) = Number of blue marbles / Total number of marbles = 900/1000 = 0.9
  • Probability of obtaining a green marble from second jar P(G) = Number of green marbles / Total number of marbles = 10/1000 = 0.01

Since these are independent events, therefore, Probability of obtaining a red, blue and a green marble from each jar = P(R)*P(B)*P(G) = 0.6*0.9*0.01 = 0.0054